# Spiking Transformers Need High Frequency Information

### v1

@@ Abstract
Spiking Transformers offer an energy-efficient alternative to conventional deep learning by transmitting information solely through binary (0/1) spikes. However, there remains a substantial performance gap compared to artificial neural networks. A common belief is that their binary and sparse activation transmission leads to information loss, thus degrading feature representation and accuracy. In this work, however, we reveal for the first time that spiking neurons preferentially propagate low-frequency information. We hypothesize that the rapid dissipation of high-frequency components is the primary cause of performance degradation. For example, on Cifar-100, adopting Avg-Pooling (low-pass) for token mixing lowers performance to 76.73%; interestingly, replacing it with Max-Pooling (high-pass) pushes the top-1 accuracy to 79.12%, surpassing the well-tuned Spikformer baseline by 0.97%. Accordingly, we introduce Max-Former that restores high-frequency signals through two frequency-enhancing operators: extra Max-Pooling in patch embedding and Depth-Wise Convolution in place of self-attention. Notably, our Max-Former (63.99 M) hits the top-1 accuracy of 82.39% on ImageNet, showing a +7.58% improvement over Spikformer with comparable model size (74.81%, 66.34 M). We hope this simple yet effective solution inspires future research to explore the distinctive nature of spiking neural networks, beyond the established practice in standard deep learning.
Figure 1: Spiking Transformer architectures: (a) Avg-Pool vs. (b) Max-Pool for token mixing, with (c) detailed implementation of the Spiking MLP (S-MLP) block. In mainstream (non-spiking) Vision Transformer research, Avg-Pool that captures global low-frequency patterns is a more common token mixing strategy than Max-Pool (high-pass) [1, 2]. Surprisingly, in Spiking Transformers, replacing Avg-Pool with Max-Pool yields a +2.39% improvement on Cifar-100.
@@ 1 Introduction
Spiking neural networks (SNNs) offer an energy-efficient alternative to conventional artificial neural networks (ANNs) [3]. Their efficiency stems primarily from spiking neurons that utilize spatiotemporal dynamics to mimic the operational principles of human brains [4]. In ANNs, all neurons within the same layer must await the complete processing of real-valued, dense tensors before any information can flow to the subsequent layer. In contrast, SNNs transmit information asynchronously, with spiking neurons consuming energy only when receiving or emitting spikes (“1”), otherwise remaining inactive [5, 6]. This binary activation pattern enables SNNs to replace the energy-intensive multiply-and-accumulate (MAC) operations that are essential in ANNs with much simpler spike-based accumulation. Leveraging the energy-efficiency benefits, Spiking Transformers, which integrate the powerful Transformer architecture with spike-based computing, have gained growing attention [7, 8].
Figure 2: Comparison between ReLU and spiking neuron (S-Neuron): (a) Input images; (b) Fourier spectrum analysis of output features processed as $\rightarrow$ $\rightarrow$ weighting, with high-frequency regions marked (red dashed boxes: regions >0.55$\times$ max amplitude) and (c) the corresponding relative log amplitude; (d) GradCAM comparison with identical architectural setting following [8], with the converted Spiking Transformer using 256 timesteps. Spiking neurons cause the rapid dissipation of high-frequency components, which consequently leads to the degradation of feature representations.
@@
However, the discrete nature of spike-based computation presents both opportunities and challenges. Currently, the most significant obstacle facing SNNs is the considerable performance gap compared to ANNs. For this discrepancy, the prevailing explanations have centered on the “representation error” [9, 10, 11], a fundamental constraint in which spiking neurons transmit information via binary spike trains rather than continuous activation values. This seems inconsistent with the established consensus in the standard deep learning literature, where networks with low-bit or even binary activations can maintain comparable performance [12, 13]. Further, it should be noted that SNNs operate on temporal sequences: despite spiking neurons strictly transmitting binary signals at each individual time step, a train of spikes spanning $n$ simulation timesteps can encode activation values with at least $\log(n)$-bit precision [14, 15]. The conflicting observation suggests unexplored opportunities for unlocking Spiking Transformers’ performance limit.
@@
It is natural to think from the frequency domain. Spiking neurons produce discrete, pulse-like activations, fundamentally distinguishing their frequency response from continuous activation functions commonly used in standard networks (e.g., ReLU [16]). Prior works have suggested that spiking neurons may enrich signals with local details (high frequencies) [17, 18]. However, in Figure 2 (b-c), we observed an intriguing phenomenon: when examining the end-to-end information flow of input $\rightarrow$ activation $\rightarrow$ weighting, rather than focusing solely on the property of activation functions, spiking neurons tend to propagate low-frequency information more prominently than ReLU.
@@
With this insight, we hypothesize that the rapid dissipation of high-frequency components is the primary cause of feature degradation, preventing Spiking Transformers from learning local, detailed information (Figure 2 (d)). To preliminarily validate this hypothesis, we conduct a simple experiment only using the non-parametric pooling operators as shown in Figure 1. From the frequency domain, Max-Pooling excels at capturing local high-frequency details (e.g., local edges/textures), whereas average pooling favors global low-frequency patterns. Notably, Spiking Transformers exhibit opposite preferences to ANNs in token mixing. While average pooling is more common for token mixing in standard Transformers [1, 2], in Spiking Transformers, replacing it with Max-Pooling yields a +2.39% improvement on CIFAR-100.
@@
In this work, we provide both theoretical and empirical evidence that high-frequency information is essential for Spiking Transformers:
  • To identify the cause of feature degradation, we reveal and theoretically prove for the first time that spiking neurons function as low-pass filters at the network level.
@@
  • We propose Max-Former to validate our hypothesis, which restores high-frequency information through two lightweight operators: extra Max-Pooling in patch embedding and Depth-Wise Convolution (DWC) in place of self-attention.
@@
  • Restoring high-frequency information significantly improves performance while saving energy cost. On ImageNet, Max-Former achieves 82.39% top-1 accuracy (+7.58% over Spikformer) with 30% energy savings at similar model size (63.99M vs. 66.34M).
@@ 2.1 Spiking Neuron Models
SNNs implement spike-driven processing through biologically-inspired neuron models for non-linear activations. The Leaky Integrate-and-Fire (LIF) model is a widely adopted abstraction of this behavior, offering an effective balance between biological plausibility and computational efficiency [19]. The discretized LIF model under each simulation timestep $n$ can be formulated as:
@@
$V[n]$ =$=\begin{cases}U[n]-V_{\mathrm{th}},&S[n]=1,\\ U[n],&S[n]=0,\end{cases}$ (3)
@@
2.2 Spiking Transformers
Drawing inspiration from biological neurons, SNNs extend conventional ANNs by incorporating temporal dynamics and discrete spike-based communication [4]. Neuromorphic chips that support such a spike-driven computing paradigm substitute power-consuming matrix-vector multiplications with sparse routing and addition operations [6, 5]. This facilitates high parallelism, scalability, and exceptional power efficiency, as they only consume tens to hundreds of milliwatts [20].
Recently, the development of Spiking Transformer has demonstrated both attractive performance and reduced energy consumption [7, 8]. Spikformer [7] pioneered the spike-based self-attention mechanism called Spiking Self Attention (SSA), which utilizes sparse spike-form Query, Key, and Value vectors, eliminating the need for energy-intensive softmax operations. Following its success, many works endeavor to enhance Spiking Transformers by adapting advanced ANN Transformer architectures [21, 22] or devising complicated spike coding mechanisms to reduce representation error (e.g., multi-threshold [23]/ multi-spike neurons [24, 25]). Here, we instead address a fundamental question: What truly limits Spiking Transformers’ performance compared to ANNs? Our investigation reveals that the answer lies in frequency properties -— specifically, that spiking neurons function as low-pass filters. In addition to theoretical proof, we probe the importance of high-frequency information through our Max-Former, which features two frequency-enhancing operators: Max-Pooling in patch embedding and DWC in place of self-attention.
@@ 3 Methods
In this section, we first present a theoretical analysis of the frequency properties of spiking neurons, , proving they act as low-pass filters at the network level despite spike trains being all-pass. This is a fundamental problem that has been overlooked in previous works. Building on this insight, we probe the importance of high-frequency information in Spiking Transformers through Max-Former, which strategically employs high-pass operators (Max-Pooling and DWC) to avoid feature degradation.
@@ 3.1 Spiking Neurons are Low-pass Filters
We begin with an intuitive time-frequency analysis using an input $x(t)=\frac{1}{3}(\sin(2\pi\cdot 100t)+\sin(2\pi\cdot 200t)+\sin(2\pi\cdot 300t))$ as shown in Figure 3. The results reveal three key observations: (1) In the time domain, the ReLU output ($r(t)$) perfectly follows $x(t)>0$ while spiking neurons selectively respond to 100Hz (Figure 3 (a)); (2) Thus, although the spiking outputs themselves, i.e., |S(f)| in the figure, exhibit slower high-frequency decay, this effect arises purely from their discrete output waveform (Figure 3 (a-b)); (3) When considering the whole process from input $\rightarrow$ activation $\rightarrow$ linear transform, ReLU expands the frequency bandwidth of $x(t)$ [26], whereas the spiking neuron exhibits strong high-frequency attenuation (Figure 3 (c)).
To further analyze the frequency-selective properties of spiking neurons, we first examine their charging process. Given Equ. (3) and Equ. (4), this can be formulated as:
@@
Equ. (8) is exactly the form of a first-order infinite-impulse-response (IIR) low-pass filter with a single pole at $z=\beta$. Accordingly, as $\beta$, i.e., the pole, approaches 1, LIF neurons exhibit stronger low-frequency selectivity.
[…]
@@
When output spike trains are weighted by a causal synaptic kernel $w[n]$, the output current $y[n]=w[n]*s[n]$ corresponds to Z–transform $Y(z)=W(z)\,S(z)$. Combining this with (8) and (10) yields the overall input-to-output transfer function
@@
Even though the synaptic kernel $W(z)$ or the spike coding process $S(z)$ changes the gain or phase response, the first-order low-pass IIR characteristics of $H(z)$ make the system $Y(z)$ inherently favor low-frequency signal components. When extending the single-layer transfer function $H^{\prime}(z)$ to $L$ layers, the low-pass term $(\frac{1-\beta}{1-\beta z^{-1}})^{L}$ further amplifies the system’s frequency selectivity. The complete formula is as follows:
@@
In the special case of the non-leaky IF neuron that obeys the charging process in (5), the resulting $H(z)$ is:
@@
Figure 4: (a) Overview of Max-Former: we restore high-frequency signals by using lightweight DWCs instead of self-attention in the early stages. Following the hierarchical design of  [27], Max-Former adopts a 3-stage architecture. $D_{i}$: feature dimensions of stage-$i$. (b) In Max-Former’s patch embedding stage, we propose three configurations (Embed-orig, Embed-Max, and Embed-Max+) to enhance high-frequency components.
@@ 3.2 Max-Former
A fundamental yet understudied opportunity in Spiking Transformers lies in their inherent low-frequency preference: while spiking neurons theoretically operate as all-pass filters, their interaction with network dynamics (e.g., membrane potential decay, threshold firing, and linear weighting) creates a tunable low-pass filtering effect. We systematically investigate this phenomenon through Max-Former. To decouple frequency effects from model complexity, we: (1) replace self-attention with high-frequency-preserving DWC in the early stages, and (2) add Max-Pooling in patch embedding to compensate spiking neurons’ low-pass preference. Notably, compared to the quadratic computational complexity of self-attention, DWC and Max-Pooling only require linear complexity relative to the sequence length and are more parameter efficient. Our work consistently adopts the LIF neuron.

### v6

# Spiking Neural Networks Need High-Frequency Information

@@ Abstract
Spiking Neural Networks promise brain-inspired and energy-efficient computation by transmitting information through binary (0/1) spikes. Yet, their performance still lags behind that of artificial neural networks, often assumed to result from information loss caused by sparse and binary activations. In this work, we challenge this long-standing assumption and reveal a previously overlooked frequency bias: spiking neurons inherently suppress high-frequency components and preferentially propagate low-frequency information. This frequency-domain imbalance, we argue, is the root cause of degraded feature representation in SNNs. Empirically, on Spiking Transformers, adopting Avg-Pooling (low-pass) for token mixing lowers performance to 76.73% on Cifar-100, whereas replacing it with Max-Pool (high-pass) pushes the top-1 accuracy to 79.12%. Accordingly, we introduce Max-Former that restores high-frequency signals through two frequency-enhancing operators: (1) extra Max-Pool in patch embedding, and (2) Depth-Wise Convolution in place of self-attention. Notably, Max-Former attains 82.39% top-1 accuracy on ImageNet using only 63.99M parameters, surpassing Spikformer (74.81%, 66.34M) by +7.58%. Extending our insight beyond transformers, our Max-ResNet-18 achieves state-of-the-art performance on convolution-based benchmarks: 97.17% on CIFAR-10 and 83.06% on CIFAR-100. We hope this simple yet effective solution inspires future research to explore the distinctive nature of spiking neural networks. Code is available: https://github.com/bic-L/MaxFormer.
Figure 1: Spiking Transformer architectures: (a) Avg-Pool vs. (b) Max-Pool for token mixing, with (c) detailed implementation of the Spiking MLP (S-MLP) block. In mainstream (non-spiking) Vision Transformer research, Avg-Pool that captures global low-frequency patterns is a more common token mixing strategy than Max-Pool (high-pass) [yu2022metaformer, yu2023metaformer]. Surprisingly, in Spiking Transformers, replacing Avg-Pool with Max-Pool yields a +2.39% improvement on Cifar-100.
@@ 1 Introduction
Spiking neural networks (SNNs) are emerging as an energy-efficient alternative to conventional artificial neural networks (ANNs) [lecun2015deep, he2024diffusion]. Their efficiency arises from spiking neurons that utilize spatiotemporal dynamics to mimic biological computation in the human brain [roy2019towards]. In ANNs, all neurons within the same layer must await the complete processing of real-valued, dense tensors before any information can flow to the subsequent layer. SNNs, however, transmit information asynchronously, with spiking neurons consuming energy only when receiving or emitting spikes (“1”), otherwise remaining inactive [akopyanTruenorthDesignTool2015, davies2018loihi]. This binary activation pattern enables SNNs to replace the energy-intensive multiply-and-accumulate (MAC) operations that are essential in ANNs with much simpler spike-based accumulation. Leveraging the energy-efficiency benefits, modern SNN variants, such as Spiking Transformers that integrate the powerful Transformer architecture with spike-based computing, have gained growing attention [zhou2022spikformer, wang2023masked].
[…]
@@
Despite their energy efficiency, the discrete nature of spike-based computation presents both opportunities and challenges. A major obstacle for SNNs remains their performance gap relative to ANNs. This gap is often attributed to the so‑called “representation error” [liDifferentiableSpikeRethinking2021, guo2022loss, dingOptimalAnnsnnConversion2021], which argues binary spike trains inherently limit the precision of feature representations compared to continuous activations. However, this seems inconsistent with the established consensus in the standard deep learning literature, where low-bit and even binary networks can still achieve comparable accuracy [li2022q, rastegari2016xnor]. Further, it should be noted that SNNs operate on temporal sequences: while spiking neurons strictly transmit binary signals at each individual time step, a train of spikes spanning $n$ simulation timesteps can encode activation values with at least $\log(n)$-bit precision [adrian1926impulses, liFreeLunchANN2021]. The conflicting observation reveals an unexplored dimension in understanding SNN performance limitations.
@@
It is natural to think from the frequency domain. Spiking neurons produce discrete, pulse-like activations, fundamentally distinguishing their frequency response from continuous activation functions commonly used in standard networks (e.g., ReLU [lecun1998gradient]). Prior works have suggested that spiking neurons may enrich signals with local details (high frequencies) [rombouts2010fractionally, fang2024spiking]. However, in Figure 2 (b-c), we observed a surprising phenomenon: when examining the end-to-end information flow of input $\rightarrow$ activation $\rightarrow$ weighting, rather than focusing solely on the property of activation functions, spiking neurons tend to propagate low-frequency information more prominently than ReLU.
@@
Feature degradation observed in SNNs may instead originate from the rapid dissipation of high-frequency components, which prevents the network from effectively capturing local, fine-grained information (Figure 2 (d)). To support this finding, we perform a simple experiment in which non‑parametric pooling operators, i.e., Max-Pool and Avg-Pool, serve as token mixers in Spiking Transformers, shown in Figure 1. From the frequency domain perspective, Max-Pool excels at capturing local high-frequency details (e.g., local edges/textures), whereas Avg-Pool favors global low-frequency patterns. Intriguing, Spiking Transformers exhibit opposite preferences to ANNs in token mixing: while standard Transformers typically employ Avg-Pool for token mixing [yu2022metaformer, yu2023metaformer], replacing it with Max-Pool in Spiking Transformers yields a +2.39% improvement on CIFAR-100, making it surpass the well-tuned Spikformer [zhou2022spikformer] baseline by 0.97%.
@@
Overall, this work provides further theoretical and empirical evidence supporting the view that high‑frequency information is essential for SNNs:
  • We provide the first theoretical proof that spiking neurons inherently act as low‑pass filters at the network level, revealing their tendency to suppress high‑frequency features.
@@
  • We propose Max-Former, which restores high‑frequency information in Spiking Transformers via two lightweight modules: extra Max-Pool in patch embedding and Depth-Wise Convolution (DWC) in place of early-stage self-attention.
@@
  • Restoring high-frequency information significantly improves performance while saving energy cost. On ImageNet, Max-Former achieves 82.39% top-1 accuracy (+7.58% over Spikformer) with 30% energy consumption and lower parameter count (63.99M vs. 66.34M).
  • Extending the insight beyond transformers, Max-ResNet-18 achieves state-of-the-art performance on convolution-based benchmarks: 97.17% on CIFAR-10 and 83.06% on CIFAR-100.
@@ 2.1 Spiking Neuron Models
SNNs implement spike-driven processing through biologically-inspired neuron models for non-linear activations. The Leaky Integrate-and-Fire (LIF) model is a widely adopted abstraction of this behavior, offering an effective balance between biological plausibility and computational efficiency [maass1997networks]. The discretized LIF model under each simulation timestep $n$ can be formulated as:
@@
$V[n]$ $=\begin{cases}U[n]-V_{\mathrm{th}},&S[n]=1,\\ U[n],&S[n]=0,\end{cases}$ (3)
@@
2.2 Spiking Neural Networks
Drawing inspiration from biological neurons, SNNs extend conventional ANNs by incorporating temporal dynamics and discrete spike-based communication [roy2019towards]. Leveraging this spike-driven mechanism, neuromorphic chips implement computation through event-driven spike routing and accumulation, which substitutes energy-intensive matrix–vector multiplications [davies2018loihi, akopyanTruenorthDesignTool2015]. This facilitates high parallelism, scalability, and exceptional power efficiency, with power consumption typically in the range of tens to hundreds of milliwatts [basu2022spiking].
Recently, the development of modern SNNs, e.g., Spiking Transformer, has demonstrated both attractive performance and reduced energy consumption [zhou2022spikformer, wang2023masked, fang2025dynamic]. Spikformer [zhou2022spikformer] pioneered the spike-based self-attention mechanism called Spiking Self Attention (SSA) that utilizes sparse spike-form Query, Key, and Value vectors to eliminate the need for energy-intensive softmax operations. Following its success, many works endeavor to enhance Spiking Transformers by adapting advanced ANN Transformer architectures [yao2024spike, zhou2024qkformer] or devising complicated spike coding mechanisms to reduce representation error (e.g., multi-threshold [hao2024lm]/ multi-spike neurons [wang2025adaptive, yao2025scaling]).
Here, we instead address a fundamental question: What truly limits SNNs’ performance compared to ANNs? Our investigation reveals that the answer lies in frequency properties -— specifically, that spiking neurons function as low-pass filters, impeding the propagation of high-frequency detail within the network. In addition to theoretical proof, we probe the importance of high-frequency information through our Max-Former, which features two frequency-enhancing operators: Max-Pool in patch embedding and DWC in place of early-stage self-attention. We further validate this principle in convolutional architectures with our proposed Max‑ResNet.
@@ 3 Methods
In this section, we first present a theoretical analysis of the frequency properties of spiking neurons. We show that, although the raw output spike trains of spiking neurons appear spectrally all‑pass due to their impulse‑shaped spike waveform, the resulting high‑frequency components are merely superficial and cannot be propagated through the network. In fact, spiking neurons act as low‑pass filters at the network level. This is a fundamental problem that has been overlooked in previous works. Building on this insight, we probe the importance of high-frequency information in SNNs through Max-Former, which strategically employs high-pass operators (Max-Pool and DWC) to restore high-frequency details and avoid feature degradation.
@@ 3.1 Spiking Neurons are Low-pass Filters
We begin with an intuitive time-frequency analysis using an input $x(t)=\frac{1}{3}(\sin(2\pi\cdot 100t)+\sin(2\pi\cdot 200t)+\sin(2\pi\cdot 300t))$ as shown in Figure 3. The results reveal three key observations: (1) In the time domain, the ReLU output ($r(t)$) perfectly follows $x(t)>0$ while spiking neurons selectively respond to 100Hz (Figure 3 (a)); (2) However, the spiking outputs’ spectral response |S(f)| still appears nearly all‑pass, which contradicts the low‑frequency behavior observed in the time domain. These spurious high‑frequency components actually arise from the impulse‑shaped spike waveform itself rather than from genuine (Figure 3 (a-b)); (3) The waveform‑induced high‑frequency components cannot be propagated across layers, resulting in a network‑level low‑pass behavior. When considering the whole process from input $\rightarrow$ activation $\rightarrow$ linear transform, ReLU expands the frequency bandwidth of $x(t)$ [kechris2024dc], whereas the spiking neuron exhibits strong high-frequency attenuation (Figure 3 (c)).
We first examine the charging process of spiking neurons to theoretically analyze their frequency-selective properties. Given Equ. (3) and Equ. (4), this can be formulated as:
@@
Equ. (8) is exactly the form of a first‐order infinite‐impulse‐response (IIR) low‐pass filter with a single pole at $z=\beta$. Accordingly, as $\beta$, i.e., the pole, approaches 1, LIF neurons exhibit stronger low‐frequency selectivity. Notably, the decay factor $\beta$ correlates to the membrane time constant $\tau$ as $\beta=1-\tfrac{1}{\tau}$, with $\tau$ ranges from $1$ to $\infty$, a smaller $\tau$ yields a smaller $\beta$. From an intuitive standpoint, a shorter time constant allows the membrane potential to respond within narrower temporal windows, rendering the neuron more sensitive to higher‑frequency inputs.
From the individual neuron to the network-level information transmission, the average membrane potential is positively and consistently correlated with the possibility of spike firing during operation. We approximate the inherently nonlinear spike‐generation process as linear around the firing threshold $V_{th}$. Denoting the firing rate by $\mathit{fr}(V)$, we define the local gain $k$ and approximate the Z–domain spike train $S(z)$ by:
$k\;=\;\left.\frac{\partial\,\mathit{fr}}{\partial V}\right|_{V=V_{th}}$ (9)
@@
When output spike trains are weighted by a causal synaptic kernel $w[n]$, the Z-transformed output current (y[n]=w[n]s[n]y[n]=w[n]*s[n]) can be denoted as Y(z)=W(z)S(z)Y(z)=W(z)\,S(z). The overall input‐to‐output transfer function is obtained by combining this with (8) and (10):
@@
The first-order low-pass IIR characteristics of $H(z)$ make the system $Y(z)$ inherently favor low-frequency signal components, regardless of whether the synaptic kernel $W(z)$ or the spike coding process $S(z)$ changes the gain or phase response. The low-pass term $(\frac{1-\beta}{1-\beta z^{-1}})^{L}$ further amplifies the system’s frequency selectivity when the process $H^{\prime}(z)$ is cascaded $L$ times (layers). The complete formula is as follows:
@@
In the special case of the non‐leaky IF neurons, which obey the charging process in (5), $H(z)$ is formulated as:
@@ 3.2 Max-Former
It remains unclear whether high‑frequency information is truly important for SNNs and whether restoring it can improve performance. Therefore, we systematically investigate the low-pass filtering characteristics of spiking neurons through Max-Former. To decouple frequency effects from model complexity, we: (1) replace self-attention with high-frequency-preserving DWC in the early stages, and (2) add Max-Pool in patch embedding to compensate spiking neurons’ low-pass preference. Notably, compared to the quadratic computational complexity of self-attention, DWC and Max-Pool only require linear complexity relative to the sequence length and are more parameter-efficient. We consistently adopt the LIF neuron model throughout this work.
Figure 4: (a) Overview of Max-Former: we restore high-frequency signals by using lightweight DWCs instead of self-attention in the early stages. Following the hierarchical design of  [liuSwinTransformerHierarchical2021], Max-Former adopts a 3-stage architecture. $D_{i}$: feature dimensions of stage-ii. (b) In Max-Former’s patch embedding stage, we propose three configurations (Embed-orig, Embed-Max, and Embed-Max+) to enhance high-frequency components.